OpenStudy (anonymous):

y = -5 when x = 3 How do I find the constant variation and then the value of y when x = -5 ?

OpenStudy (amistre64):

youd have to know how x and y relate to each other first

OpenStudy (amistre64):

that is why we need a constant of variation; and how they vary; directly, jointly, etc ...

OpenStudy (anonymous):

Explain please?

OpenStudy (amistre64):

without being told the relationship between y and x; there is no good grounds with which to determine a "constant of variation". spose they are inversely related; then 5 = k/3 ; such that k = 15 y = 15/x ; when x = -5, y = -3 if they are directly related, then: 5 = k3; and k = 5/3 y = 5x/3 ; when x=-5, y = -25/3 there is no good way to determine the relationship between y and x in order to solve the problem with any degree of confidence

OpenStudy (amistre64):

y and x could have a linear relationship such that: y = x + 2 5 = 3 + 2 y = -5 + 3 = -2

OpenStudy (anonymous):

So what would be the constant variation of y = -2 when = 2 ? I have to figure out if it's a direct variation first right ?

OpenStudy (amistre64):

you would definantly have to figure out what the relationship between x and y is first; yes